OpenStudy (anonymous):

HELP PLLEEAASS ON A TIME LIMIT WILL MEDAL AND FAN PLLLLLLLLZZZZZZZZ Solve 6 1/3+K+3 5/6+5 1/2=26 1/6

OpenStudy (anonymous):

HELP

OpenStudy (88):

1. Rename the mixed numbers with common denominators 2. Add the mixed numbers together to get 15 4/6 3. Subtract the sum from 26 to get 10 1/2 4. Since the only term left on the left is k so k=10 1/2

OpenStudy (anonymous):

so u have to add all of the fractions then subtract what you got from 26 1/6 to get K

OpenStudy (anonymous):

so do I change 6 1/3 to 19/3 and then change all the ether ones the same way

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and when you do that find the lcd for all of them

OpenStudy (anonymous):

do I also change the answer of the question not for K but the one after the = sine

OpenStudy (anonymous):

here i will give you the fractions

OpenStudy (anonymous):

okay since ur denominators are 3, 6, and 2 u have to find a the smallest number that can be divided by 3, 6, 2 any guesses?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

or 2

OpenStudy (anonymous):

3, 6, and 2 dont go into 1 :) 1 goes into them a number that 3, 6, 2 go into is 12

OpenStudy (anonymous):

o sorry 12

OpenStudy (anonymous):

no its ok ur trying :)

OpenStudy (anonymous):

ok so put the number 12 aside and turn the mixed numbers into improper fractions do you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so 6 1/3 is going to be what?

OpenStudy (anonymous):

19/3

OpenStudy (anonymous):

good now what is 3 5/6 doing to be?

OpenStudy (anonymous):

23/6

OpenStudy (anonymous):

grat and 5 1/2?

OpenStudy (anonymous):

11/2

OpenStudy (anonymous):

lol great*

OpenStudy (anonymous):

okay so u know have 19/3, 23/2, and 11/2

OpenStudy (anonymous):

now you have to find out how many times the denominator goes into 12 how many times does 3 go into 12?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

great so you multiply 3*4 to get 12 right? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so what you do to the bottom you do to the top and since you multiplied the bottom by 4 you multiply the top by 4, so what is 19*4?

OpenStudy (anonymous):

yes witch=76

OpenStudy (anonymous):

great so you have 76/12 move on to the next fraction 23/6 how many times does 6 go into 12?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

okay so you would multiply the numerator by 2 so what is 23*2?

OpenStudy (anonymous):

46

OpenStudy (anonymous):

oh great! :) so since you have 46 put it over 12, and your fraction is going to be 46/12

OpenStudy (anonymous):

so can you do 11/2 by yourself or do you want my help wit it still? :)

OpenStudy (anonymous):

I can do it just how do I find out K by doing this and the answer to 11/2 is 23/2

OpenStudy (anonymous):

not quite 2 goes into 12 6 times so you would multiply 11*6 and get 66 then put that over 12 to get 66/12

OpenStudy (anonymous):

o ya sorry i new that just my mind went blank

OpenStudy (anonymous):

no its ok so then u add 76/12+46/12+66/12 but dont add the denominator just the numerators then add /12 to the answer you get :)

OpenStudy (anonymous):

188 is 76+46+66

OpenStudy (anonymous):

great so 188/12

OpenStudy (anonymous):

ok turn 26 1/6 into an improper fraction

OpenStudy (anonymous):

157/6 i think

OpenStudy (anonymous):

great so how many times does 6 go into 12?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so then u multiply the top by 2 so 156*2=?

OpenStudy (anonymous):

312

OpenStudy (anonymous):

great put that over 12 so u have 312/12 and subtract 188/12 remember since the denominators are the same leave them alone so just do 312-188 and put it over /12

OpenStudy (anonymous):

124/12

OpenStudy (anonymous):

wonderful! now turn that into a mixed number do you know how to do that?

OpenStudy (anonymous):

yes it is 10 4/12 witch also = 2/6 witch also = 1/3 so 10 1/3

OpenStudy (anonymous):

YES! wonderful and 10 1/3 is what K is!

OpenStudy (anonymous):

ok thanks sooooooo much can you help with one more

OpenStudy (anonymous):

Y/9=4 2/3

OpenStudy (anonymous):

ok ill try i gtg soon tho

OpenStudy (anonymous):

ok since there equal fractions turn 4 2/3 into a improper fraction

OpenStudy (anonymous):

k i gtg just turn that into an improper fraction multiply the numerator by 3 since 3*3= 9 and whatever you get =Y